A solution is made by mixing 45.0 mL of ethanol, C2H6O, and 55.0 mL of water. Assuming ideal behavior, what is the vapor pressure of the solution at 20 °C?

Constants @ 20*C:

ethanol = 0.789(Density) & 43.9(Vapor Pressure - Torr)

water = 0.998(Density) & 17.5(Vapor Pressure - Torr)
Answer #1

P = PaXa + PbXb

mass of ethanol = 45 x 0.789= 35.5 gm = 35.5/46 = 0.772 moles,

mass of water = 0.998 x 55 = 54.89 gm = 54.89/18 = 3.05 moles ,

Xa = (0.772)/(0.772+3.05) = 0.202 , Xb = 1-0.202 = 0.798,

P = (0.202)(43.9) + (0.798)(17.5) = 22.83 Torr